ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 190, 2012

Baumgartner, U., R. G. Möller and G. A. Willis Hyperbolic groups have flat-rank at most 1

365

Bell, J. P. Transcendence degree of division algebras  195
Besser, A. On the syntomic regulator for K_{1} of a surface  29
Blaier, N. A characterization of the disc via a Hessian equality  1

Brech, C. and P. Koszmider On universal Banach spaces of density continuum

93

Caprace P.-E. and N. Monod A lattice in more than two Kac– Moody groups is arithmetic

413

Cheong, O., X. Goaoc and A. Holmsen Lower bounds to Helly numbers of line transversals to disjoint congruent balls

213

Culler, M. and P. B. Shalen Margulis numbers for Haken manifolds

445

De la Rosa, M., L. Frerick, S. Grivaux and A. Peris Frequent hypercyclicity, chaos, and unconditional Schauder decompositions

389

Drmota, M. and J. F. Morgenbesser Generalized Thue–Morse sequences of squares

157

Einsiedler, M. and S. Kadyrov Entropy and escape of mass for

SL_{3}(ℤ)\ SL_{3}(ℝ) 253

Ferreira, W. and P. Roitman Area preserving transformations in two-dimensional space forms and classical differential geometry

325

Frerick, L. See M. De la Rosa Glasner, E. and M. Megrelishvili On fixed point theorems and

nonsensitivity

289

Goaoc, X. See O. Cheong

Griesmer, J. T. Sumsets of dense sets and sparse sets  229

Grivaux, S. See M. De la Rosa Holmsen, A. See O. Cheong Kadyrov, S. See M. Einsiedler Kang, J. Existence of leafwise intersection points in the unrestricted

case

111

Koszmider, P. See C. Brech

Kowalski, P. Schanuel property for additive power series  349

Kupferman, R. and Y. Shamai Incompatible elasticity and the immersion of non-flat Riemannian manifolds in Euclidean space

135

Megrelishvili, M. See E. Glasner Möller , R. G. See U. Baumgartner Monod, N. See Caprace P.-E. Morgenbesser, J. F. See M. Drmota Munteanu, R.-B. A non-product type non-singular transformation

which satisfies Krieger’s property A

307

Newelski, L. Model theoretic aspects of the Ellis semigroup  477

Peris, A. See M. De la Rosa Pinasco, D. and I. Zalduendo A probabilistic approach to polyno-

mial inequalities

67

Qian, G. A character theoretic criterion for a p-closed group  401

Roitman, P. See W. Ferreira Shalen, P. B. See M. Culler Shamai, Y. See R. Kupferman Spielman, D. A. and N. Srivastava An elementary proof of the

restricted invertibility theorem

83

Srivastava, N. See D. A. Spielman Willis, G. A. See U. Baumgartner Zalduendo, I. See D. Pinasco